OpenStudy (anonymous):

find the product 2 1/8*1 1/6*3

OpenStudy (anonymous):

Please help

OpenStudy (austinl):

I don't mean to be rude or anything, but I must say; the way you have this displayed has me kinda confused.

OpenStudy (austinl):

\(\displaystyle \bf 2\frac{1}{8}\cdot1\frac{1}{6}\cdot3\) Is this what it looks like?